# pnasresearcharticle

### v1

@@ Conclusion
We have performed measurements of the dissipation of 3He fourth-sound resonances for several different confinements. We find that the temperature scaling of the dissipation is well-described by a vortex mutual friction model for a fixed vortex density. The lack of power dependence suggests that the density of vortices is set in upon phase transition rather than being extrinsically generated by flow (see discussion in Supplementary Materials (8)). We find that the density of vortices is independent of temperature ramp rate, but is a function of the channel thickness. A modified version of the Kibble-Zurek model is proposed in which a spatial length scale of the system sets the estimated defect density rather than the quench rate as is the case in the Kibble-Zurek model. This interpretation is found to be within an order of magnitude of the observations, comparable to the agreement seen in tests of the Kibble-Zurek mechanism, and the scaling with confinement is in reasonable agreement. We expect that similar finite-size effects may be present in a variety of systems where the average domain size set by the freeze-out time is small compared to one or more dimensions of the system.
[…]
The authors thank C.B. Winkelmann for generously providing code used in the computation of the effective viscosity of superfluid 3He, W. Zurek for extensive discussions, and E. Varga for nanofabrication assistance. The authors note that the pre-print (27) is highly relevant to this work, but appeared shortly after the submission of this manuscript. This work was supported by the University of Alberta, Faculty of Science and Office of the Vice President of Research; the Natural Sciences and Engineering Research Council, Canada (Grants No. RGPIN-2022-03078 and No. CREATE-495446-17); and the Canada Foundation for Innovation.
[…]

### v2

@@ Conclusion
We have performed measurements of the dissipation of 3He fourth-sound resonances for several different confinements. We find that the temperature scaling of the dissipation is well-described by a vortex mutual friction model for a fixed vortex density. The lack of power dependence suggests that the density of vortices is set in upon phase transition rather than being extrinsically generated by flow (see discussion in Supplementary Materials [8]). We find that the density of vortices is independent of temperature ramp rate, but is a function of the channel thickness. A modified version of the Kibble-Zurek model is proposed in which a spatial length scale of the system sets the estimated defect density rather than the quench rate as is the case in the Kibble-Zurek model. This interpretation is found to be within an order of magnitude of the observations, comparable to the agreement seen in tests of the Kibble-Zurek mechanism, and the scaling with confinement is in reasonable agreement. We believe this result demonstrates the non-trivial dependence of defect density on the boundaries of the system. This may be relevant to a variety of systems where the Kibble-Zurek length scale is small compared to a single dimension of the system.